Find a particular solution to the differential equation using the Method of Undetermined Coefficients. 6t x(t) – 12x (t) +
Answer #1

x - 12x +36x= 3tect mt Let x=en be a trial solution of x- 12x + 36x=0. Then the auxiliary equation is m²-12 m+36=0 m - 6)